A 2.5kg mass tied to a string is whirled around a horizontal circle of radius 1.7m. The tension in the string is 35.7N and the string is inclined at 60° above the horizontal. Calculate the speed at which the stone moves round the circle.


1
Expert's answer
2021-01-23T10:29:29-0500

According to the second Newton's law "F=ma". So, we have


"N\cos60°=m\frac{v^2}{r} \to v=\sqrt{\frac{N\cos60°\cdot r}{m}}="


"=\sqrt{\frac{35.7\cdot\cos60°\cdot 1.7}{2.5}}\approx3.5(m/s)" . Answer